Classifying Balance Sheet and Income Statement Accounts and Computing Current Ratio
Shoprite Holdings Ltd is an African food retailer listed on the Johannesburg Stock Exchange. The following accounts are selected from its annual report for the fiscal year ended June 30, 2014. The amounts below are in millions of South African Rand.

a. Indicate the appropriate classification of each account as appearing in either its balance sheet (B) or its income statement (I).

(Rand millions) Amount Classification
Sales of merchandise R 107,254
Depreciation and amortisation 2,154
Reserves (Retained earnings) 14,036
Property, plant & equipment 14,000
Cost of goods and services 91,494
Trade and other payables 17,150
Total assets 45,583
Total equity 18,101
Employee benefits expense 8,797
Total non-current assets 16,548
Total non-current liabilities 6,349


b. Using the data, compute Shoprite's total liabilities at June 30 2014.
R Answer million

c. Calculate Shoprite's current ratio as of June 30, 2014.
Round to two decimal places.
Answer
Answer #1

Answer

  • [a]

(Rand millions)

Amount

Classification

Sales of merchandise

R 107,254

Income Statement (I)

Depreciation and amortisation

R 2,154.00

Income Statement (I)

Reserves (Retained earnings)

R 14,036.00

Balance Sheet (B)

Property, plant & equipment

R 14,000.00

Balance Sheet (B)

Cost of goods and services

R 91,494.00

Income Statement (I)

Trade and other payables

R 17,150.00

Balance Sheet (B)

Total assets

R 45,583.00

Balance Sheet (B)

Total equity

R 18,101.00

Balance Sheet (B)

Employee benefits expense

R 8,797.00

Income Statement (I)

Total non-current assets

R 16,548.00

Balance Sheet (B)

Total non-current liabilities

R 6,349.00

Balance Sheet (B)

[b]

A

Total assets

R 45,583.00

B

Total equity

R 18,101.00

C = A - B

Total Liabilities

R 27,482.00

[c]

A

Total assets

R 45,583.00

B

Total non-current assets

R 16,548.00

C = A - B

Total Current Assets

R 29,035.00

D

Total Liabilities

R 27,482.00

E

Total non-current liabilities

R 6,349.00

F = D - E

Total Current Liabilities

R 21,133.00

G = C/F

Current Ratio

1.37